# NAC-TCN: Temporal Convolutional Networks with Causal Dilated Neighborhood Attention for Emotion Understanding

### 12 Dec 2023 v1

@@ Abstract.
ccs: General and reference Experimentationccs: General and reference Performanceccs: Human-centered computing Collaborative and social computing devicesccs: Computer systems organization Neural networksccs: Computing methodologies Scene understandingccs: Computing methodologies Vision for roboticsccs: Computing methodologies Activity recognition and understandingccs: Computing methodologies Computer vision tasksccs: Computing methodologies Computer visionccs: Computing methodologies Computer vision problemsccs: Computing methodologies Machine learning approaches
@@ 4.2. Implementation Details
We use an Adam Optimizer with a base learning of $0.001$ alongside an annealing cosine scheduler. We use a batch size of 16 for the Affwild2, EmoReact, and AFEW-VA datasets and all models were trained for 10 epochs. AFEW and Affwild all used subject based k-fold cross validation to ensure that information leakage did not occur between testing and training. The validation dataset of AffWild2 was used as the evaluation set and kept separate from training data. The EmoReact dataset had preset train and test splits that were used to be in line with the performance of prior models. The best performing model was selected for each method. The same random seed value was selected to ensure reproducibility. The number of heads for DiNAT was selected using hyperparameter search ($\{2^{n}\mid 1\leq n\leq 3\}$).

### 06 Jan 2024 v2

@@ Abstract.
journalyear: 2023copyright: acmlicensedconference: 2023 the 7th International Conference on Video and Image Processing (ICVIP); December 14–17, 2023; Kyoto, Japanbooktitle: 2023 the 7th International Conference on Video and Image Processing (ICVIP) (ICVIP 2023), December 14–17, 2023, Kyoto, Japandoi: 10.1145/3639390.3639392isbn: 979-8-4007-0938-8/23/12ccs: General and reference Experimentationccs: General and reference Performanceccs: Human-centered computing Collaborative and social computing devicesccs: Computer systems organization Neural networksccs: Computing methodologies Scene understandingccs: Computing methodologies Vision for roboticsccs: Computing methodologies Activity recognition and understandingccs: Computing methodologies Computer vision tasksccs: Computing methodologies Computer visionccs: Computing methodologies Computer vision problemsccs: Computing methodologies Machine learning approaches
@@ 4.2. Implementation Details
We use an Adam Optimizer with a base learning of $0.001$ alongside an annealing cosine scheduler. We use a batch size of 16 for the AffWild2, EmoReact, and AFEW-VA datasets and all models were trained for 10 epochs. AFEW and AffWild2 all used subject based k-fold cross validation to ensure that information leakage did not occur between testing and training. The validation dataset of AffWild2 was used as the evaluation set and kept separate from training data. The EmoReact dataset had preset train and test splits that were used to be in line with the performance of prior models. The best performing model was selected for each method. The same random seed value was selected to ensure reproducibility. The number of heads for DiNAT was selected using hyperparameter search ($\{2^{n}\mid 1\leq n\leq 3\}$). Note that for AffWild2, we build off the open source library provided by Nguyen 22 for testing our new model to ensure consistency when comparing to other methods (Nguyen_2022_CVPR, ).